OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 2|7-7x| = 2x+4

OpenStudy (anonymous):

divide be 2 first to get the absolute by itself

OpenStudy (anonymous):

What do I divide? the 7 by 2?

OpenStudy (anonymous):

divide the left side by 2 and teh right side by 2. This will get rid of the 2 on the left and simplify the right

OpenStudy (anonymous):

\[\left| 7-7x \right|=x+2\]

OpenStudy (anonymous):

And then do I cancel the x's out?

OpenStudy (anonymous):

7-7x is the agruement and it can equal +(x+2) or -(x+2)

OpenStudy (anonymous):

not yet..

OpenStudy (anonymous):

Do you understand my statement just now

OpenStudy (anonymous):

I kinda do, a little bit is still confusing

OpenStudy (anonymous):

what is inside the absolute value symbol whether positive or negative comes out positive right?

OpenStudy (anonymous):

so the expression is saying that whatever 7-7x is it can be the positive or negative of x+2

OpenStudy (anonymous):

http://www.themathpage.com/alg/absolute-value.htm

OpenStudy (anonymous):

So I take 7 & divide it by 2? Sorry I'm a little lost

OpenStudy (anonymous):

you need to set 7-7x=x+2 and solve for x as well as set 7-7x=-x-2 and solve for x

OpenStudy (anonymous):

This gets you both of your solutions

OpenStudy (anonymous):

treat the absolute value signs as ( )

OpenStudy (anonymous):

Would it be 3/2 & 5/8?

OpenStudy (anonymous):

\[\frac{ 2\left| 7-7x \right| }{ 2 }=\frac{ 2x+4 }{ 2 }\]

OpenStudy (anonymous):

see how the 2 cancels from the left and the right becomes x+2??

OpenStudy (anonymous):

I havent solved it yet, hold on

OpenStudy (anonymous):

ya I got the same

OpenStudy (anonymous):

Alright I think I got the hang of it now thanks! (=

OpenStudy (anonymous):

yup